# Edge Weight Prediction For Category-Agnostic Pose Estimation

### v1

@@ 3.1 Edge Weight Prediction
The architecture of our skeleton prediction network is illustrated in Figure 4. Our first step is to refine the keypoints features to enhance their structural understanding. We find the graph-decoder proposed in [14] effective for this purpose, as it is designed to consider structural dependencies. We observe that further refinement of the input image features, in conjunction with the keypoint features, can significantly improve structure learning. To achieve this, we add a cross-attention mechanism, where the input image features are updated by attending to the keypoint features. We refer to this module as a dual-attention graph decoder, which encourages the exchange of structural information between the image and keypoint features. Once the feature refinement is complete, we apply a multi-headed self-attention mechanism to predict the residual adjacency matrix $\Delta A$.
To ensure a smooth combination of the prior and residual matrices, while minimizing disruptions in early training, we incorporate the zero-convolution operator used in ControlNet [53]. The resulting combined matrix is then passed through an activation function to ensure positivity:
$A^{\prime}=ReLU(A_{prior}+c_{i}\Delta A)$ (2)
where $c_{i}$ is a learnable scalar initialized to zero, so the output skeleton begins as the prior adjacency matrix $A_{prior}$, stabilizing the model early in training. As training progresses, the model gradually learns to balance the contributions of $A_{prior}$ and $\Delta A$.
@@
Lastly, we enforce adjacency matrix symmetry and normalization to ensure proper interpretation as a valid skeleton, and right-stochasticity. Symmetry is enforced by averaging the matrix with its transpose and normalization is applied row-wise: $A=\frac{A^{\prime}+A^{\prime T}}{2}$ (3) $\tilde{A}_{ij}=\frac{A_{ij}}{\sum_{j}A_{ij}}$ (4)
Figure 4: Skeleton Predictor Model. Our skeleton predictor contains a feature refinement module that uses the prior input graph $A_{prior}$ and the support image features $F_{s}$ to enhance the keypoint features $F_{s}^{k}$. Self-attention is then applied to predict the residual adjacency output $\Delta A$.
3.2 Adjacency Supervision
Optimizing the skeleton predictor using only localization loss leads to sub-optimal results. Thus, we seek to add to the predicted adjacency matrix a dedicated supervision signal. However, even for humans, identifying the correct graph structure is not always straightforward, and determining appropriate edge weights is particularly challenging. Given this uncertainty, an unsupervised approach is necessary to learn these relations effectively. SDPNet [31] suggested using a GCN with a masking strategy to supervise the skeleton prediction. We advocate this strategy, as we aim for an adjacency matrix that can help the model overcome occlusions using structure. However, this requires additional GCN training which is not used during inference. Therefore, some of the structure data may be encoded in the GCN weights and not only in the adjacency matrix.
@@ 3.2 Adjacency Supervision
Thus, we adopt the masking strategy, but instead of using an auxiliary GCN, we use the keypoint prediction module for supervision. Optimization is done as follows: We first randomly mask a portion of the input support keypoint features using the vector mask $M$, and replace them with a learnable masking token $F_{mask}$. Then, using the masked input, we predict the query keypoint locations:
$P_{i}^{m}=g_{\theta}(M\circ F_{s}+(1-M)\circ F_{mask},\tilde{A},P^{0})$ (5)
@@
where $\circ$ is the Hadamard product, $g_{\theta}$ is the graph-decoder of the keypoint prediction module, $P^{0}$ are the predicted coordinates from the proposal generator, and $\tilde{A}$ is the predicted adjacency matrix. The adjacency loss is the $L_{1}$ distance between the predicted keypoints using masked features and the ground-truth locations $\hat{P}$:
@@
We want to encode this data into the adjacency matrix. Therefore, during the backward pass for this loss, we freeze the decoder’s weights and all the inputs except $\tilde{A}$, to optimize only the adjacency predictor. In addition, we follow [19] and drop the heatmap loss used in GraphCape [14] as we empirically notice that this loss term doesn’t improve performance. Thus, training is done using:
@@
where $\mathcal{L}_{offset}$ is the $L_{1}$ localization loss as in [14]. More details about our training scheme are in the supplementary.
Support GT CapeFromer [36] SCAPE [19] GraphCape [14] EdgeCape
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 5: Qualitative Comparison. We visualize keypoint predictions for the 1-shot setting. The left column shows the support data, followed by ground-truth query keypoints, and results from different methods.
3.3 Feature Extraction
Recent visual foundation models like Dino [6, 26] have demonstrated that their self-supervised pre-trained features can serve as powerful descriptors for semantic correspondence [51]. Trained on extensive datasets, these models are highly suited for category-agnostic applications, as they capture generalizable features that are not bound to specific categories. Recent efforts have focused on identifying features best suited for semantic correspondence [2, 52]. We aim to leverage these analyses for the task of CAPE (the differences between the two tasks are discussed in [36, 19]). However, GraphCape [14] showed that DinoV2 backbone performed worse than SwinV2, indicating that Swin backbone is better suited for the suggested architecture (aligning with Lin et al. [20]).
Nonetheless, we advocate the use of Dino features, as Banani et al. [10] demonstrated that Dinov2 possesses an intrinsic 3D structural awareness. This structural awareness is beneficial for capturing skeletal structures — the focus of our work. As Dino models may capture distinct properties across various network layers, relying solely on the final layer’s features may not be optimal for localization tasks. Following [10], we adopt a dense, multi-scale probing approach similar to the DPT decoder [29] to harness the nuanced features at different scales. By freezing the pre-trained backbone and training only the DPT layers, we preserve the generalization capabilities of self-supervised features, which are crucial for category-agnostic tasks. This is compared to previous methods, which fine-tuned the pre-trained backbone.
3.4 Markovian Attention Bias
In category-agnostic pose estimation (CAPE), encoding spatial relations between keypoints enhances robust localization, especially when dealing with novel objects [14, 33]. While Transformer models provide a global receptive field, this flexibility lacks structural constraints, requiring explicit encoding of these relations.
We build on GraphCape [14], which integrates GCN layers to propagate structural information, but these are limited to nearest-neighbor connections. To capture more complex and distant dependencies, we incorporate a bias term in the Transformer’s self-attention mechanism based on graph connectivity [48]. To support weighted edges that represent varying strengths in keypoint relations, we adopt a continuous formulation. We follow Ma et al. [21] and treat the normalized adjacency matrix $\widetilde{A}$ as a stochastic matrix representing a Markov process, where $(\widetilde{A}^{k})_{ij}$ indicates the probability of transitioning from node $v_{i}$ to $v_{j}$ in $k$-hops. This enables both direct and multi-hop relations to influence attention. We start by constructing the matrix $P$:
where $K$ is the maximum hops considered. Then, we use the following self-attention term:
$a_{ij}=\frac{(h_{i}W_{Q})(h_{j}W_{K})^{T}}{\sqrt{d}}+\text{MLP}(P_{ij})$ (9)
where the MLP: $\mathbb{R}^{K}\to\mathbb{R}$ modulates the influence of keypoints based on graph distance. The full formulation and derivation of this bias term are in the supplementary.
4 Experiments
Model Backbone 1-Shot 5-Shot
POMNet [42] ResNet-50 84.23 78.25 78.17 78.68 79.17 79.70 84.72 79.61 78.00 80.38 80.85 80.71
CapeFormer [35] ResNet-50 89.45 84.88 83.59 83.53 85.09 85.31 91.94 88.92 89.40 88.01 88.25 89.30
ESCAPE [24] ResNet-50 86.89 82.55 81.25 81.72 81.32 82.74 91.41 87.43 85.33 87.27 86.76 87.63
MetaPoint+ [7] ResNet-50 90.43 85.59 84.52 84.34 85.96 86.17 92.58 89.63 89.98 88.70 89.20 90.02
SDPNet  [31] HRNet-32 91.54 86.72 85.49 85.77 87.26 87.36 93.68 90.23 89.67 89.08 89.46 90.42
X-Pose [44] Swin 89.07 85.05 85.26 85.52 85.79 86.14 - - - - - -
PPM  [27] Stable Diffusion 91.03 88.06 84.48 86.73 87.40 87.54 93.64 92.71 91.76 92.85 91.94 92.58
SCAPE [^1] Evaluated without keypoint identifiers, which were shown to be inapplicable for real-world scenarios [24, 44, 14]  [19] DinoV2 91.47 86.29 87.23 87.07 86.94 87.80 94.33 90.53 91.49 90.68 89.80 91.37
GraphCape [14] SwinV2 91.19 87.81 85.68 85.87 85.61 87.23 94.24 91.32 90.15 90.37 89.73 91.16
Ours DinoV2 + DPT 93.69 89.27 87.85 86.67 87.59 89.01 95.51 91.94 91.33 90.36 91.92 92.21
Table 1: MP-100 Results. PCK performance under 1-shot and 5-shot settings. Our approach outperforms other methods on the 1-shot setting. In addition, on the 5-shot setting, we outperform similar-sized methods and reduce the performance gap with PPM [27], a method that uses a huge pre-trained diffusion model and test-time optimization. The best results are bold, and the second-best are underlined.
In line with prior CAPE works, we use the MP-100 dataset for training and evaluation. This dataset consists of over 20,000 images drawn from various category-specific pose estimation datasets, covering 100 object categories. Each category includes varying numbers of keypoints, with up to 68 keypoints per category. The dataset is divided into five distinct, mutually exclusive splits, ensuring that categories used for training, validation, and testing do not overlap.
To evaluate our model’s performance, we use the Probability of Correct Keypoint (PCK) [46] metric, following established practices. We set the PCK threshold to 0.2, as commonly done in previous works. In the supplementary, we comprehensively compare our method by including two additional metrics: Area Under the ROC Curve (AUC) and Normalized Mean Error (NME).
4.1 Implementation Details
For a fair comparison, training parameters, data augmentations, and data pre-processing are kept the same as in previous works. In addition, we use the smallest version of DinoV2 to match the backbone size of GraphCape [14]. More implementation details are in the supplementary.
@@ 4.2 Qualitative Results
Figure 5 provides a qualitative comparison between our method and previous CAPE works, including CapeFormer [36], SCAPE [19], and GraphCape [14]. Our method’s incorporation of predicted weighted pose graphs acts as a stronger prior for keypoint localization, effectively breaking symmetry and enhancing localization accuracy.
Additionally, we provide a few examples of graph predictions. Figure 6 shows input prior graphs $A_{prior}$ and the predicted refined graphs, which serve as inputs to the graph-based keypoint prediction module. As can be seen, our network weakens symmetric parts’ connections, which can hurt localization. For example, the table has stronger connections between the base and legs than between the different corners of the base. Alternatively, it creates new helpful connections in the human face. Finally, in Figure 7 we show PCA visualizations of our feature extraction method compared to previous methods. Our approach captures fine-grained details and learns asymmetrical features that enhance localization.
In the supplementary, we also demonstrate the instance adaptability of our graph prediction module by presenting graph examples from different instances within the same category. Additionally, we provide additional predicted graphs across various categories, more comparison examples, and PCA visualizations of features from different backbones.
Input SwinV2 DinoV2 DinoV2 +
DPT Probing
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Feature Extraction. PCA visualization comparing different feature extraction models. Our approach captures fine-grained details and asymmetrical features that aid localization.
@@ 4.3 Quantitative Results
We report results on the MP-100 dataset under 1-shot and 5-shot settings in Table 1. For a fair comparison, we show results using similar-sized backbones. It is worth noting, however, that PPM [27] differs from other methods as it is based on the significantly larger Stable Diffusion model [32]. Moreover, it requires test-time optimization, which is unnecessary for the other methods. Nevertheless, we include their results for completeness. As can be seen, our method outperforms the state-of-the-art by an average of 1.21% under the 1-shot setting, and achieves the highest performance among similar-sized methods under the 5-shot setting, with a margin of 0.84%.

### v2

@@ 3.2 Category-Agnostic Pose-Graph Prediction
where $A_{prior}\in\{0,1\}^{K\times K}$ is the unweighted graph input provided by the user (like in GraphCape), and $F_{s}\in\mathbb{R}^{hw\times C}$ and $F_{s}^{k}\in\mathbb{R}^{K\times C}$ are the support image and keypoint features.
Refining Structural Features.
First, we refine the keypoint features $F_{s}^{k}$ to enhance their structural semantics. Rather than sharing information only among the keypoint features $F_{s}^{k}$, we also incorporate the support image features $F_{s}$ to provide valuable global structural context, which is especially important in the category-agnostic setting where objects’ orientations vary widely and keypoints alone may be ambiguous.
We build on a graph transformer decoder (Hirschorn and Avidan, 2024), which is well-suited for modeling structural dependencies. To strengthen the interaction between $F_{s}^{k}$ and $F_{s}$, we add a cross-attention layer where the image features are updated using the keypoint features. We refer to this as a dual-attention graph decoder, since it enables bidirectional exchange of structural information between image and keypoint features. In practice, we find that refining image features alongside keypoint features significantly improves the quality of learned structural representations. The architecture of our pose-graph prediction network is illustrated in the Supplementary (Figure 15).
This modification also changes the role of the decoder itself. Originally designed to exchange information between support keypoints and a query image, it now jointly processes keypoints and support image features of the same image. This repurposing allows the model to learn richer structural priors directly from the support image, leveraging the fact that structural patterns are shared within object categories.
Predicting Residual Edges.
Using the refined structure-aware features $F_{refined}^{k}$, we compute $\Delta A$ as pairwise cosine-similarities:
$\Delta A_{ij}=<F_{refined}^{i},F_{refined}^{j}>$ (2)
This similarity measure naturally quantifies the directional alignment and strength of relations between keypoints. We found this similarity measure as a balance between accuracy and efficiency. Alternatives such as MLPs or attention-based predictions yielded only marginal improvements at a higher computational cost.
The resulting $\Delta A$ is then combined with the user-provided prior $A_{\text{prior}}$. A naïve approach would directly sum $A_{\text{prior}}$ and $\Delta A$, but this leads to unstable training and poor convergence, as the output graphs at the start of training lack structural meaning. Instead, we introduce a simple yet effective scaling mechanism that stabilizes training. We combine $A_{\text{prior}}$ and $\Delta A$ using a learnable scaling factor initialized to zero, and pass the result through an activation function to ensure positivity:
$A^{\prime}=\text{ReLU}(A_{\text{prior}}+c\Delta A)$ (3)
where $c$ is the learnable scalar. At the start of training, $c=0$, meaning the output pose graph is initially $A_{\text{prior}}$, providing a stable foundation.
@@
Lastly, we enforce adjacency matrix symmetry and normalization to ensure proper interpretation as a valid pose-graph, and right-stochasticity. Symmetry is enforced by averaging the matrix with its transpose and normalization is applied row-wise: $A=\frac{A^{\prime}+A^{\prime T}}{2}$ (4) $\tilde{A}_{ij}=\frac{A_{ij}}{\sum_{j}A_{ij}}$ (5)
Pose-Graph Prediction Self-Supervision.
Training the pose-graph predictor with only a localization loss does not reliably produce useful graphs. We therefore introduce an additional supervision signal. Designing such a signal is challenging: even for humans, identifying the correct graph structure is non-trivial, and assigning precise edge weights is impractical. To address this, we adopt an unsupervised masking strategy, where our keypoint prediction module provides indirect supervision for the pose-graph predictor. Using masking, we aim to find adjacency matrices that can help the model overcome occlusions, a realistic scenario for this task.
@@
Optimization is done as follows. We randomly mask a subset of the input support keypoint features using the vector mask $M$, and replace them with a learnable masking token $F_{mask}$. Then, using the masked input, the model predicts the query keypoint locations:
$P_{i}^{m}=g_{\theta}(M\circ F_{s}^{k}+(1-M)\circ F_{mask},\tilde{A},P^{0})$ (6)
@@
where $\circ$ is the Hadamard product, $g_{\theta}$ is the keypoints prediction module, $P^{0}$ are the predicted coordinates from the proposal generator (Section C), and $\tilde{A}$ is the predicted adjacency matrix. The adjacency loss is the $L_{1}$ distance between the predicted locations $P^{m}$ using masked keypoint features and the ground-truth locations $\hat{P}$:
@@
During backpropagation, this loss is used to update only the adjacency predictor. To achieve this, we freeze the decoder’s weights and all inputs except $\tilde{A}$, ensuring that the adjacency matrix itself encodes information useful for overcoming occlusion.
The final training objective combines this adjacency loss with the standard localization loss:
@@
where $\mathcal{L}_{offset}$ is $L_{1}$ localization loss as in (Shi et al., 2023b). More details about our training scheme are in the supplementary, Section C.4.
3.3 Markov Attention Bias
In CAPE, encoding spatial relations between keypoints is critical for robust localization (Hirschorn and Avidan, 2024; Rusanovsky et al., 2024). Having established a way to predict more informative adjacency matrices, we now focus on how to better leverage these spatial relations during localization. Keypoint localization is done using a transformer decoder (Figure 3) where the self-attention mechanism provides a global receptive field by considering all pairwise keypoint interactions. However, this flexibility comes at a cost, as it weakens the influence of structural relations. We aim to improve the model’s ability to capture both local and long-range structural dependencies. GraphCape (Hirschorn and Avidan, 2024) partially tackled this issue by integrating GCNs into the feed-forward network. While effective at modeling local structure, GCNs still struggle with reasoning over long-range relations and therefore do not provide a complete solution. To address this limitation, we introduce a new graph-informed bias term (Ying et al., 2021) into the self-attention mechanism, enabling structural connectivity patterns to directly influence the flow of information and enhance overall performance.
Our predicted adjacency matrix $\widetilde{A}$ contains weighted edges, reflecting varying strengths of keypoint relations. We normalize it row-wise, treating it as a stochastic matrix that defines a Markov process. Under this view, $(\widetilde{A}^{k})_{ij}$ represents the probability of reaching node $v_{j}$ from node $v_{i}$ in exactly $k$-hops. $\widetilde{A}_{ij}$ captures direct (1-hop) links while $(\widetilde{A}^{k})_{ij}$ captures multi-hop (long-range) relations, providing a principled way to encode both local and global structure. We start by constructing the matrix $P$:
where $K$ is the maximum number of hops considered. Intuitively, $P_{ij}$ captures multi-hop structural relations. Then, using an MLP: $\mathbb{R}^{K}\to\mathbb{R}$, we modulate the influence of keypoints based on graph distance, allowing the model to learn which relations are most relevant for localization. The resulting modified self-attention in the decoder is:
$a_{ij}=\frac{(h_{i}W_{Q})(h_{j}W_{K})^{T}}{\sqrt{d}}+\text{MLP}(P_{ij})$ (10)
This formulation preserves the expressiveness of self-attention while explicitly guiding it to respect structural priors. In effect, the model learns to weigh keypoint interactions based on both semantic similarity and graph connectivity. A figure of the new graph-transformer decoder, along with the full derivation of the bias term and training scheme, are provided in the supplementary, Section C.3.
4 Experiments
We follow standard CAPE practice by training and evaluating on MP-100 dataset (Xu et al., 2022a), which includes over 20,000 images from 100 categories with up to 68 keypoints per category. The dataset is split such that categories used for training, validation, and testing do not overlap, and it provides five different category splits to ensure robust generalization. We also report results on AP-10K (Yu et al., 2021) (supplementary, Section B.2), an animal pose dataset covering 23 families and 60 species. Unlike MP-100, it uses a shared keypoint definition and pose-graph structure.
For evaluation, we use the Probability of Correct Keypoint (PCK) (Yang and Ramanan, 2012) metric. We use a 0.2 threshold for backward compatibility with prior works. But, as this metric became saturated, we follow (Chen et al., 2025a) and also report mPCK (mean PCK over thresholds 0.05, 0.1, 0.15, 0.2) for a more challenging evaluation.
Implementation Details.
For a fair comparison, we match the training parameters, data augmentations, and data pre-processing to previous works. Additional details are in the supplementary (Section C.5). Notably, our method introduces only minimal latency overhead, adding a lightweight MLP (Markovian Attention Bias) and a single decoder pass (pose-graph predictor), adding only $\sim$2 ms per image (on an A5000 GPU). A detailed analysis of computation and complexity is in the supplementary.
Support GT CapeFromer GraphCape Ours
@@ 4 Experiments
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 4: Qualitative Comparison. We visualize keypoint predictions for the 1-shot setting. The left column shows the support data, followed by ground-truth query keypoints, and results of different methods. Our method performs best by leveraging predicted weighted pose-graphs, which serve as more effective structural priors for keypoint localization.
@@ 4.1 Qualitative Results
Figure 4 provides a qualitative comparison between our method, CapeFormer (Shi et al., 2023b), and previous graph-based CAPE work GraphCape (Hirschorn and Avidan, 2024). CapeFormer does not use any graph prior, while GraphCape uses pre-defined graph prior. With stronger pose-graphs and a better mechanism to exploit them via Markov Attention Bias, our approach achieves superior localization.
Additionally, Figure 5 provides examples of category-agnostic pose-graph predictions. As can be seen, our network weakens symmetric parts’ connections, which can hurt localization. For example, the table has stronger connections between the base and legs than between the different corners of the base. Alternatively, it creates new helpful connections in the human face. In the supplementary (Section B.3), we demonstrate the instance adaptability of our graph prediction module.
@@
Figure 5: Predicted Pose-Graphs. We visualize predicted graphs: left column shows input $A_{\text{prior}}$, right shows output AA. Line width reflects edge weight. Observe the slimmer table-base edges and the new facial edges. Our model prunes symmetric part links and forms connections that aid localization.
@@ 4.2 Quantitative Results
Following (Chen et al., 2025a), we categorize CAPE methods into two groups: graph and point-based methods. We note that point-based methods largely pursue orthogonal directions and can be combined. Additional metrics and in-depth comparisons are in the supplementary (Section B.1).
Table 1: MP-100 Results. We compare our method against baselines without graph priors or using fixed graphs. Our method consistently outperforms others across all settings and data splits.
Method Graph Prior 1-Shot 5-Shot
CapeFormer None 78.44 73.56 73.61 73.14 73.51 74.45 83.71 79.77 79.18 79.62 75.89 79.63
GraphCape Fixed 79.87 75.06 76.16 73.58 73.98 75.73 83.93 79.81 78.78 79.02 79.28 80.16
Ours Predicted 81.96 77.63 77.35 75.68 75.98 77.72 84.92 80.91 79.68 79.76 80.08 81.07
Graph-based methods. Table 1 reports mPCK results of graph-based methods on MP-100 under 1-shot and 5-shot settings. CapeFormer serves as a baseline, not using any prior, while GraphCape uses fixed pose-graphs. As shown, our method consistently outperforms other methods across all splits and settings. We achieve significant gains of +1.99% in 1-shot and +0.9% in 5-shot mPCK over GraphCape. At a finer evaluation threshold (e.g., 0.05), these margins widen substantially, reaching +4.0% and +2.23% respectively. Importantly, the improvements are stable across splits - even in cases where GraphCape underperforms CapeFormer (for example, split 3 in 5-shot), our approach remains superior. This suggests that simply injecting graph priors can be brittle, whereas predicting them yields a stronger and more reliable structural prior. Overall, we reach a substantial improvement of 3.27% (1-shot) and 1.44% (5-shot) mPCK over CapeFormer.
Table 2: Point-based Results. PCK@0.2 performance comparison. Our approach outperforms other methods on the 1-shot and 5-shot settings. Best results are bold, second-best are underlined.
Model 1-Shot 5-Shot
POMNet (Xu et al., 2022a) 84.23 78.25 78.17 78.68 79.17 79.70 84.72 79.61 78.00 80.38 80.85 80.71
ESCAPE (Nguyen et al., 2024) 86.89 82.55 81.25 81.72 81.32 82.74 91.41 87.43 85.33 87.27 86.76 87.63
MetaPoint+ (Chen et al., 2024) 90.43 85.59 84.52 84.34 85.96 86.17 92.58 89.63 89.98 88.70 89.20 90.02
SDPNet  (Ren et al., 2024) 91.54 86.72 85.49 85.77 87.26 87.36 93.68 90.23 89.67 89.08 89.46 90.42
X-Pose (Yang et al., 2024) 89.07 85.05 85.26 85.52 85.79 86.14 - - - - - -
SCAPE  (Liang et al., 2024) 91.47 86.29 87.23 87.07 86.94 87.80 94.33 90.53 91.49 90.68 89.80 91.37
FMMP (Chen et al., 2025a) 88.19 85.26 83.03 85.29 84.72 85.30 91.97 89.36 87.35 89.33 88.56 89.31
CapeFormer (Shi et al., 2023a) 91.07 86.94 87.05 85.53 85.77 87.27 94.98 91.47 90.69 90.24 88.62 91.20
GraphCape (Hirschorn and Avidan, 2024) 92.39 88.47 89.24 85.76 86.66 88.50 95.21 91.51 90.65 89.92 90.22 91.50
Ours 93.57 89.53 89.31 87.38 87.29 89.42 95.45 91.88 91.13 90.28 90.61 91.87
Point-based methods. We show in Table 2 our results compared to recent CAPE models using the more common, but saturated, PCK threshold 0.2. EdgeCape surpasses the previous state-of-the-art, with margins that are on par with - and in many cases exceeds - the gains reported by prior leading methods. As shown, at PCK@0.2 the improvements over GraphCape are +0.92% (1-shot) and +0.37% (5-shot), highlighting that performance using this coarse threshold is approaching saturation. Nonetheless, our method achieves impressive gains of +1.62% (1-shot) and +0.5% (5-shot) over SCAPE, the previous non-graph-based state-of-the-art. Notably, our work can be further combined with other point-based CAPE methods, which we leave for future work.
Limitations. Our method consistently enhances accuracy across both low- and higher-shot settings, with particularly notable improvements in the 1-shot scenario. Yet, the performance gain is smaller in higher-shot settings - a trend also observed when comparing CapeFormer and GraphCape. This is expected, as graph-based methods are particularly effective for handling occlusions and preserving structure under low-shot conditions. Nonetheless, our contribution in higher-shot settings remains substantial, as our learned pose-graphs and their effective utilization enable consistently stronger gains than those achieved by GraphCape.